OpenStudy (anonymous):

What is the limit of abs(x-2)/(x-2) as x approaches 2?

OpenStudy (helder_edwin):

use one-sided limits

OpenStudy (anonymous):

abs(x-2)/(x-2) means just the absolute value of (x-2) divided by (x-2).

OpenStudy (helder_edwin):

\[ \large \lim_{x\to2}\frac{|x-2|}{x-2} \] right?

OpenStudy (anonymous):

Exactly

OpenStudy (helder_edwin):

compute separatedly \[ \Large \lim_{x\to2+}\frac{|x-2|}{x-2} \] and \[ \Large \lim_{x\to2-}\frac{|x-2|}{x-2} \] (one-sided limits)

OpenStudy (anonymous):

Alright, I understand the graphical process of one-sided limits, but is there an analytical (algebraic) approach?

OpenStudy (helder_edwin):

just follow this: \[ \Large x\to2+\Rightarrow x>2\Rightarrow x-2>0\Rightarrow |x-2|=x-2 \]

OpenStudy (dumbcow):

you could change the abs value to x-2 for x>2 2-x for x<2 take limit of each case

OpenStudy (anonymous):

Does the limit here actually exist? Since the limit from the left isn't equal to the limit from the right.

OpenStudy (helder_edwin):

no it does not

OpenStudy (anonymous):

So would the answer be "does not exist b/c..." or the two different values it approaches from either side?